The pressure 75.0 m under water is 837 kPa. What is the pressure of the atmospheres (atm)? Also, what is this pressure in millimeters of Mercury (mmHg)?
P=
P=
1
Expert's answer
2020-04-02T08:56:46-0400

 To calculate gauge pressure at height h=75 , use formula P=hdg+P∘ where P∘ is atmospheric pressure,  h (=75 m) is the height below the open water surface, d (=1 g/cm3 ) is the density of water and g (=9.8 m/s2)is the acceleration due to gravity. Then,

P∘=837 - 9.8*75*1=102 kPa ≈ 1atm ≈ 760 mmHg;

Answer: The pressure of the atmosphere is 1atm or 760 mmHg.